# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

### v1

@@ Abstract
With the rapid development of XR, 3D generation and editing are becoming more and more important, among which, stylization is an important tool of 3D appearance editing. It can achieve consistent 3D artistic stylization given a single reference style image and thus is a user-friendly editing way. However, recent NeRF-based 3D stylization methods face efficiency issues that affect the actual user experience and the implicit nature limits its ability to transfer the geometric pattern styles. Additionally, the ability for artists to exert flexible control over stylized scenes is considered highly desirable, fostering an environment conducive to creative exploration. In this paper, we introduce StylizedGS, a 3D neural style transfer framework with adaptable control over perceptual factors based on 3D Gaussian Splatting (3DGS) representation. The 3DGS brings the benefits of high efficiency. We propose a GS filter to eliminate floaters in the reconstruction which affects the stylization effects before stylization. Then the nearest neighbor-based style loss is introduced to achieve stylization by fine-tuning the geometry and color parameters of 3DGS, while a depth preservation loss with other regularizations is proposed to prevent the tampering of geometry content. Moreover, facilitated by specially designed losses, StylizedGS enables users to control color, stylized scale and regions during the stylization to possess customized capabilities. Our method can attain high-quality stylization results characterized by faithful brushstrokes and geometric consistency with flexible controls. Extensive experiments across various scenes and styles demonstrate the effectiveness and efficiency of our method concerning both stylization quality and inference FPS.
@@ 1 Introduction
Nowadays, the once professionally-dominated domain of artistic content creation has become increasingly accessible to novice users, thanks to recent groundbreaking advancements in visual artistic stylization research. As a pivotal artistic content generation tool in crafting visually engaging and memorable experiences, 3D scene stylization has attracted growing research efforts. Previous methodologies have attempted style transfer by enabling control over diverse explicit representations such as mesh [1, 2, 3], voxel [4, 5], and point cloud [6, 7, 8]. However, the quality of their results is limited by the quality of the geometric reconstructions. The recent 3D stylization methods benefit from the emerging implicit neural representations [9, 10, 11], such as neural radiance field (NeRF) [12, 13, 14, 15, 16], achieving more faithful and consistent stylization within 3D scenes. Nonetheless, NeRF-based methods are computationally intensive to optimize and suffer from the geometry artifacts in the original radiance fields.
@@
The recently introduced 3D Gaussian Splatting (3DGS) [17], showcasing remarkable 3D reconstruction quality from multi-view images with high efficiency, suggests representing the 3D scene using an array of colored and explicit 3D Gaussians. Given that practical 3D stylization applications often demand a prompt response, we propose using 3DGS as the representation of real-world scenes and conducting 3D stylization on it. Recent 3DGS scene manipulation methods [18, 19, 20, 21] explore the editing and control of 3D Gaussians using text instructions within designated regions of interest or semantic tracing. However, these approaches are constrained by text input and fall short of delivering detailed style transfer capabilities. The methods for performing 3D Gaussian stylization remain less explored in the field.
In this paper, we introduce the first controllable scene stylization method based on 3DGS, StylizedGS. With a single reference style image, our method can effectively transfer its style features to the entire 3D scene, represented by a set of 3D Gaussians. It facilitates the artistic creation of visually coherent novel views that exhibit transformed and detailed style features in a visually reasonable manner. More importantly, StylizedGS operates at a notable inference speed, ensuring the efficient synthesis of stylized scenes. To address the desire for users to control perceptual factors such as color, scale, and spatial aspects, as in 2D image style transfer [22, 23, 24] and 3D scene style transfer [25], we enhance the flexibility and introduce an advanced level of perceptual controllability in our approach to achieve personalized and diverse characteristics in 3DGS stylization.
The proposed 3DGS stylization process is formulated as an optimization problem that optimizes the geometry and color of 3D Gaussians to render images with faithful stylistic features while preserving the semantic content. A trivial solution is directly minimizing both a style loss and a content loss to align the rendered images with the style image and avoid overly strong stylization by fine-tuning the color of 3D Gaussians, respectively. Nevertheless, it poses a challenge in obtaining high-quality results with intricate style details as shown in Fig. 10 (b), as learning only the color alterations cannot effectively capture the overall style pattern. Instead, we propose to learn both the optimal geometry and color parameters of 3D Gaussians to capture the detailed style feature and facilitate the stylization of the entire 3D scene. We propose a 2-step stylization framework with the first step to enhance the fidelity in the stylized scene by the color match and reduce cloudy artifacts and geometry noises by the proposed GS filter. The second step is the optimization for stylization with a nearest neighbor feature match (NNFM) loss. We also introduce a depth preservation loss without the need for additional networks or regulation operations to preserve the overall learned 3D scene geometry. Last, we specially design a set of effective loss functions and optimization schemes to enable flexible perceptual control for users, including color, scale, and spatial regions Our contribution can be summarized as follows:
@@
  • We introduce StylizedGS, a novel controllable 3D Gaussian stylization method that organically integrates various modules with proper improvements to transfer detailed style features and produce faithful novel stylized views.
@@
  • We empower users with an efficient stylization process and flexible control by specially designed losses, enhancing their creative capabilities.
@@ 2 Related Work
Image Style Transfer. Style transfer aims to generate synthetic images with the artistic style of given images while preserving content. Initially proposed in neural style transfer methods by Gatys et al. [26, 27], this process involves iteratively optimizing the output image using Gram matrix loss and content loss calculated from VGG-Net [28] extracted features. Subsequent works [29, 30, 31, 32] have explored alternative style loss formulations to enhance semantic consistency and capture high-frequency style details such as brushstrokes. Feed-forward transfer methods [33, 34, 35], where neural networks are trained to capture style information from the style image and transfer it to the input image in a single forward pass, ensuring faster stylization. Recent improvements in style loss [31, 32, 13] involve replacing the global Gram matrix with the nearest neighbor feature matrix, improving texture preservation. Some methods adopt patch matching for image generation, like Fast PatchMatch [36] and PatchMatch [37], but are limited to specific views. Combining neural style transfer with novel view synthesis methods without considering 3D geometry may lead to issues like blurriness or view inconsistencies.
@@
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [17] has emerged as a robust approach for real-time radiance field rendering and 3D scene reconstruction. Recent methods [20, 19, 18, 38] enhance semantic understanding of 3D scenes and enabled efficient text-based editing using pre-trained 2D models. Despite these advancements, existing 3DGS works lack support for image-based 3D scene stylization that faithfully transfers detailed style features while offering flexible control. In this work, leveraging the 3DGS representation, we significantly reduce both training and rendering times for stylization, enabling interactive perceptual control over stylized scenes.
@@
Figure 2: StylizedGS Pipeline. We first reconstruct a photo-realistic 3DGS $G_{\theta}^{rec}$ from multi-view input. Following this, color matching with the style image is performed, accompanied by the application of 3DGS filters to preemptively address potential artifacts. Throughout optimization, we employ multiple loss terms to capture detailed local style structures and preserve geometric attributes. Users can flexibly control color, scale, and spatial attributes during stylization through customizable loss terms. Once this stylization is done, we can obtain consistent free-viewpoint stylized renderings.
@@
3D Style Transfer. 3D scene style transfer aims to transfer the style to the scene with both style fidelity and multi-view consistency. With the increasing demand for 3D content, neural style transfer has been expanded to various 3D representations. Stylization on meshes often utilizes differential rendering to propagate style transfer objectives from rendered images to 3D meshes, enabling geometric or texture transfer [3, 1, 2]. Other works, using point clouds as the 3D proxy, ensure 3D consistency when stylizing novel views. For instance, [7] employs featurized 3D point clouds modulated with the style image, followed by a 2D CNN renderer to generate stylized renderings. However, explicit methods’ performance is constrained by the quality of geometric reconstructions, often leading to noticeable artifacts in complex real-world scenes.
@@
Therefore implicit methods such as NeRF [39] have gained considerable attention for their enhanced capacity to represent complex scenes. Numerous NeRF-based stylization networks incorporate image style transfer losses [26, 13] during training or are supervised a mutually learned image stylization network [12] to optimize color-related parameters based on a reference style. Approaches like [14, 40] support both appearance and geometric stylization to mimic the reference style, achieving consistent results in novel-view stylization. However, these methods involve time-consuming optimization and exhibit slow rendering due to expensive random sampling in volume rendering. They also lack user-level flexible and accurate perceptual control for stylization.
@@
While controlling perceptual factors, such as color, stroke size, and spatial aspects, has been extensively explored in image domain style transfer [22, 23, 24], the application of perceptual control in 3D stylization has not been well utilized. [41, 42] establish semantic correspondence in transferring style across the entire stylized scene but only limited to spatial control and doesn’t allow users to interactively specify arbitrary regions. ARF-plus [25] introduce more perceptual controllability into the stylization of radiance fields, yet the demand for enhanced flexibility and personalized, diverse characteristics in 3D stylization remains unmet.
@@
Our approach achieves rapid stylization within a minute of training, ensuring real-time rendering capabilities. It adeptly captures distinctive details from the style image and preserve recognizable scene content with fidelity. Additionally, we empower users with perceptual control over color, scale, and spatial factors for customized stylization.
@@ 3 Method
Given a 3D scene represented by a collection of captured images with corresponding camera parameters, our objective is to efficiently accomplish consistent style transfer from a 2D style image with arbitrary style to the 3D scene. Building on recent advances in 3D Gaussian Splatting (3DGS) [17], we model the 3D scene with the 3DGS and then optimize the 3DGS representation to generate stylized scenes with intricate artistic features and high visual quality. Our stylization process consists of two key steps: 1) We first recolor the 3D scene to align its color statistics with those of the given style image. Simultaneously, a specific 3D Gaussian filter is applied to minimize the impact of floaters in the reconstruction, which is crucial for the final stylization effect. 2) Then in the optimization phase, we exploit nearest-neighbor feature matching style loss to capture detailed local style patterns. Additionally, we incorporate a depth preservation loss, and some regularization terms to preserve geometric contents. Finally, in Sec. 3.3, we introduce how to achieve flexible control over the color, scale, and spatial attributes in our 3DGS stylization. As a result, users can create stylized renderings with customized artistic expression, and explore the consistently stylized scene in a free-view manner.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting [17] encapsulates 3D scene information using a suite of 3D colored Gaussians. This technique exhibits rapid inference speeds and exceptional reconstruction quality compared to NeRF. To represent the scene, each Gaussian is described by a centroid $p=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D vector $s\in\mathbb{R}^{3}$ for scaling, and a quaternion $q\in\mathbb{R}^{4}$ for rotation. Additionally, an opacity value $\alpha\in\mathbb{R}$ and a color vector $c$ represented in the coefficients of a spherical harmonic (SH) function of degree three are used for fast alpha-blending during rendering. These trainable parameters are collectively symbolized by $G_{\theta_{i}}$, where $G_{\theta_{i}}={p_{i},s_{i},q_{i},\alpha_{i},c_{i}}$, representing the parameters for the $i$-th Gaussian. To visualize the 3D Gaussians and supervise their optimization, 3DGS projects them onto the 2D image plane. The implementation leverages differentiable rendering and gradient-based optimization on each pixel for the involved Gaussians. The pixel color $c^{\alpha}$ is determined by blending the colors $c_{i}$ of those ordered Gaussians that overlap the pixel. This process can be formulated as:
@@ 3.2 Style Transfer to 3D Gaussian Splatting
Given a set of multi-view images $\mathcal{I}_{content}$, we first obtain the reconstructed 3DGS model $G_{\theta}^{rec}$ to represent the original scene. Our goal is to transform $G_{\theta}^{rec}$ to a stylized 3DGS model $G_{\theta}^{sty}$ which matches the detailed style features of a 2D style image $\mathcal{I}_{style}$ and preserve the content of the original scene.
Color Match. To enhance the alignment of hues with the style image and achieve flexible color control in the stylized results, we initially recolor the 3DGS $G_{\theta}^{rec}$ through color transfer from the style image. Drawing inspiration from [22], we employ a linear transformation of colors in RGB space, followed by a color histogram matching procedure between the style images and training views. Let $\{p_{s}^{i}\}$ be the set of all pixels in the style image, $\{p_{c}^{i}\}$ be the set of all pixels in the content image to be recolored, then we solve the following linear transformation to align the mean and covariance of color distributions between the content image set and the style image:
$p_{c}^{re}=\mathbf{A}p_{c}+b,\,c^{re}=\mathbf{A}c+b$
$s.t.\,E[p_{c}^{re}]=E[p_{s}],Cov[p_{c}^{re}]=Cov[p_{s}]$
where $\mathbf{A}\in\mathbb{R}^{3\times 3}$ and $b\in\mathbb{R}^{3}$ are the solution to the linear transformation, $p_{c}^{re}$ is the the recolored pixels of the content images. The color parameter $c$ of 3DGS $G_{\theta}^{rec}$ is transformed to the recolored color parameter $c^{re}$. Please refer to the supplementary document for the mathematical derivation.
Consequently, this alignment process ensures that the recolored training images $\mathcal{I}_{content}^{re}$ and the color attributes of 3DGS $c^{re}$, are consistent with the color palette of the style reference. However, some floaters in the original 3DGS will also be colored during this process, significantly affecting the quality of the stylization, as shown in Fig. 11. These colored floaters often result from Gaussians with excessively large scales or improper opacity values. To address this issue, we propose applying a 3DGS filter to exclude these floaters before the stylization process. After filtering certain 3D Gaussians, we fine-tune $G_{\theta}^{rec}$ to minimize the impact on the overall scene reconstruction quality and enhance color correspondence. During fine-tuning, we incorporate the reconstruction loss from [17]:
@@
Stylization. After matching the color and filtering possible artifacts from 3DGS reconstruction, an intuitive and effective strategy for 3D scene stylization is to leverage features extracted by a pre-trained convolutional neural network (e.g., VGG [28]) and calculate a set of loss functions between training views and the style image. To transfer detailed high-frequency style features from a 2D style image to a 3D scene, we exploit the nearest neighbor feature matching concept inspired by [13, 43], where we minimize the cosine distance between the feature map of rendered images and its nearest neighbor in the style feature map. As illustrated in Fig. 2, we extract the VGG feature maps $F_{style}$, $F_{content}$, and $F_{render}$ for $\mathcal{I}_{style}$, $\mathcal{I}_{content}^{re}$, and the rendered images $\mathcal{I}_{render}$, respectively. The nearest neighbor feature match loss is formulated as:
@@
[…]
Here, $D(a,b)$ is the cosine distance between two feature vectors $a$ and $b$. To preserve the original content structure during the stylization, we additionally minimize the mean squared distance between the content feature $F_{content}$ and the stylized feature $F_{render}$:
@@
where $H$ and $W$ represent the size of rendered images. While the content loss $\mathcal{L}_{content}$ mitigates the risk of excessive stylization, optimizations applied to 3DGS geometric parameters can still induce changes in scene geometry. To address this, we introduce a depth preservation loss to maintain general geometric consistency without relying on an additional depth estimator. Therefore, there exists a trade-off between geometric optimization and depth preservation loss, enabling our stylization to effectively transfer style patterns while mitigating substantial detriment to geometric content. Specifically, in the color match stage, we initially generate the original depth map $D_{origin}$ by employing the alpha-blending method of 3DGS. The alpha-blended depth $d^{\alpha}$ of each pixel is calculated as follows:
$d^{\alpha}=\sum_{i}T_{i}\alpha_{i}d_{i}$ (6)
where $d_{i}$ is the depth value for $i$-th Gaussian. During the stylization, we minimize the $L_{2}$ loss between the rendered depth map $D_{render}$ and the original one $D_{origin}$:
@@
We also add some regularization terms that perform on the changes of scale $\Delta s$ and opacity $\Delta\alpha$ value in $G_{\theta}^{sty}$:
@@
Figure 3: Color control Results. Our approach facilitates versatile color management in stylized outputs, allowing users to either retain the scene’s original hues or apply distinct color schemes from alternative style images. Users have the autonomy to transfer the entire style, the pattern style only, or mixed arbitrary pattern and color styles.

### v2

@@ Abstract
As XR technology continues to advance rapidly, 3D generation and editing are increasingly crucial. Among these, stylization plays a key role in enhancing the appearance of 3D models. By utilizing stylization, users can achieve consistent artistic effects in 3D editing using a single reference style image, making it a user-friendly editing method. However, recent NeRF-based 3D stylization methods encounter efficiency issues that impact the user experience, and their implicit nature limits their ability to accurately transfer geometric pattern styles. Additionally, the ability for artists to apply flexible control over stylized scenes is considered highly desirable to foster an environment conducive to creative exploration. To address the above issues, we introduce StylizedGS, an efficient 3D neural style transfer framework with adaptable control over perceptual factors based on 3D Gaussian Splatting (3DGS) representation. We propose a filter-based refinement to eliminate floaters that affect the stylization effects in the scene reconstruction process. The nearest neighbor-based style loss is introduced to achieve stylization by fine-tuning the geometry and color parameters of 3DGS, while a depth preservation loss with other regularizations is proposed to prevent the tampering of geometry content. Moreover, facilitated by specially designed losses, StylizedGS enables users to control color, stylized scale, and regions during the stylization to possess customization capabilities. Our method achieves high-quality stylization results characterized by faithful brushstrokes and geometric consistency with flexible controls. Extensive experiments across various scenes and styles demonstrate the effectiveness and efficiency of our method concerning both stylization quality and inference speed.
@@ 1 Introduction
Nowadays, the once professionally-dominated domain of artistic content creation has become increasingly accessible to novice users, thanks to recent groundbreaking advancements in visual artistic stylization research. As a pivotal artistic content generation tool in crafting visually engaging and memorable experiences, 3D scene stylization has attracted growing research efforts. Previous methods have achieved attractive 3D scene style transfer results over diverse explicit representations such as mesh [1, 2, 3], voxel [4, 5], and point cloud [6, 7, 8]. However, the quality of their results is limited by the precision of the geometric reconstructions. The recent 3D stylization methods benefit from the emerging implicit neural representations [9, 10, 11], such as Neural Radiance Fields (NeRF) [12, 13, 14, 15, 16], achieving more faithful and consistent stylization within 3D scenes. Nonetheless, NeRF-based methods are computationally intensive to optimize and suffer from the geometry artifacts of the original radiance fields.
@@
The recently introduced 3D Gaussian Splatting (3DGS) [17], showcasing remarkable 3D reconstruction quality from multi-view images with high efficiency, suggests representing the 3D scene using an array of colored and explicit 3D Gaussians. Performing 3D stylization on scene represented by 3DGS brings the benefits of prompt response and flexible style control to 3D stylization applications due to the explicit nature of 3DGS. Recent 3DGS scene manipulation methods [18, 19, 20, 21] explore the editing and control of 3D Gaussians using text prompts within designated regions of interest or semantic tracing. However, these approaches are driven by text input and fall short of delivering detailed style transfer capabilities. Some styles are also difficult to describe through simple text prompt. Recently appeared studies on 3DGS stylization  [22, 23] show limited capabilities of learning and transferring style features. Notably, these works also lack comprehensive control over stylization effects.
In this paper, we introduce the first controllable scene stylization method based on 3DGS, StylizedGS. Given a reference style image, our method effectively transfer the style features to the entire 3D scene represented by a set of 3D Gaussians. Our method facilitates the artistic creation of visually coherent novel views that exhibit transformed and detailed style features in a visually reasonable manner. More importantly, StylizedGS operates at a fast inference speed, ensuring efficiency when rendering stylized scenes. To achieve effective stylization and avoid geometry distortion, a straightforward approach is to fine-tune the color of 3D Gaussians, minimizing both style loss and content loss. However, focusing solely on color fails to effectively capture the overall stylistic features when the style patterns contain intricate details. Instead, we propose formulating the 3DGS stylization as a joint optimization of both the geometry and color of 3D Gaussians to capture the detailed style features while preserving the semantic content. Specifically, our method employs a two-stage stylization framework. In the first stage, we optimize the 3D Gaussians to align the color distribution with the style image while reducing artifacts and geometry noises using our filter-based 3D Gaussian refinement. In the second stage, we achieve final stylization by applying a nearest neighbor feature match (NNFM) loss, optimizing the 3D Gaussians’ parameters to capture coherent fine-grained style details.
To enable users to customize the stylization process, we propose a series of strategies and loss functions to control perceptual factors such as color, scale, and spatial regions  [24, 25, 26, 27] in the final stylized results. By incorporating this enhanced level of perceptual controllability, our approach synthesizes diverse 3D scene stylization results with varied visual characteristics. Furthermore, we introduce a depth preservation loss that eliminates the necessity for additional networks or regularization operations to preserve the learned 3D scene geometry. It establishes a delicate balance between geometric optimization and depth preservation, facilitating effective style pattern transfer while mitigating significant deterioration of geometric content.
Our contribution can be summarized as follows:
@@
  • We introduce StylizedGS, a novel controllable 3D Gaussian stylization method that organically integrates the filter-based 3DGS refinement and the depth preservation loss with other stylization losses to transfer detailed style features and produce faithful novel stylized views.
@@
  • We empower users with an efficient stylization process and flexible control through specially designed learning schemes and losses, enhancing their creative capabilities.
@@ 2 Related Work
Image Style Transfer. Style transfer aims to generate synthetic images with the artistic style of given images while preserving content. Initially proposed in neural style transfer methods [28, 29], this process involves iteratively optimizing the output image using Gram matrix loss and content loss calculated from VGG-Net [30] extracted features. Subsequent works [31, 32, 33, 34] have explored alternative style loss formulations to enhance semantic consistency and capture high-frequency style details such as brushstrokes. Feed-forward transfer methods [35, 36, 37], where neural networks are trained to capture style information from the style image and transfer it to the input image in a single forward pass, ensuring fast stylization. Recent improvements in style loss [33, 34, 13] involve replacing the global Gram matrix with the nearest neighbor feature matrix, improving texture preservation. Some methods adopt patch matching for image generation, like PatchMatch [38] and Fast PatchMatch [39], but can only be applied to limited views. Recent works [40, 41] explore popular tools such as CycleGAN [42], Transformer [43], and diffusion model [44] for style transfer. However, directly applying these 2D methods to 3D stylization will lead to issues like blurriness or view inconsistencies due to the inadequate consideration of 3D geometry.
@@
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [17] has emerged as an approach for real-time radiance field rendering and 3D scene reconstruction. Recent methods [20, 19, 18, 45] enhance semantic understanding of 3D scenes and enable efficient text-based editing using pre-trained 2D models. Additionally, geometry deformation [46] and texture editing [47] have also been explored on 3DGS. Please refer to the survey [48, 49] for more details. Despite these advancements, existing 3DGS works lack support for image-based 3D scene stylization that faithfully transfers detailed style features while offering flexible control.
@@
Figure 2: StylizedGS Pipeline. We first reconstruct a photo-realistic 3DGS $G_{\theta}^{rec}$ from multi-view input. Following this, color matching with the style image is performed, accompanied by the filter-based refinement to preemptively address potential artifacts. During optimization, we employ multiple loss terms to capture detailed local style structures and preserve geometric attributes. Users can flexibly control color, scale, and spatial attributes during stylization through customizable loss terms. Once this stylization is done, we can obtain consistent free-viewpoint stylized renderings.
@@
3D Style Transfer. With the increasing demand for 3D content, neural style transfer has been expanded to various 3D representations. Stylization on meshes often utilizes differential rendering to transfer style from rendered images to 3D meshes, enabling geometric or texture transfer [3, 1, 2]. Other works, using point clouds as the 3D proxy, ensure 3D consistency when stylizing novel views. LSNV [7] employs featurized 3D point clouds modulated with the style image, followed by a 2D CNN renderer to generate stylized renderings. However, explicit methods’ performance is constrained by the quality of geometric reconstructions, often leading to noticeable artifacts in complex real-world scenes.
@@
Implicit methods, such as NeRF [50], have gained considerable attention for their enhanced capacity to represent complex scenes. Many NeRF-based stylization works incorporate image style transfer losses [28, 13] during training or adopt a mutually learned image stylization network [12, 10] to optimize color-related parameters given a reference style image. Approaches like [14, 51] support both appearance and geometric stylization to mimic the reference style, achieving consistent results in novel-view stylization. However, these methods involve time-consuming optimization and exhibit slow rendering due to expensive random sampling in volume rendering. They also lack user-level flexible and accurate perceptual control for stylization.
@@
While controlling perceptual factors, such as color, stroke size, and spatial aspects, have been extensively explored in image style transfer [24, 25, 26], they remain underdeveloped for 3D stylization. [52, 53] establish semantic correspondence in transferring style across the entire stylized scene, but they primarily focus on spatial control and lack the capability for users to interactively specify arbitrary regions. ARF-plus [27] introduces more perceptual controllability into the stylization of radiance fields, yet the demand for enhanced flexibility and personalized, diverse characteristics in 3D stylization remains unmet.
@@
In this work, leveraging the 3DGS representation, we achieve rapid stylization within a minute of training, ensuring real-time rendering capabilities. Recent arXiv works [22, 23] on 3DGS-based style transfer exhibit limitations in maintaining intricate style details and coherence. [22] is feedforward-based and struggles to capture locally coherent style features and patterns with AdaIN [54], while [23] simply adopts classic style loss to optimize the color of Gaussians. Our method not only effectively captures distinctive details from the style image and preserves recognizable scene content with fidelity, but also empowers users with perceptual control over color, scale, and spatial factors for customized stylization.
@@ 3 Method
Given a 3D scene represented by a collection of captured images with corresponding camera parameters, our objective is to efficiently accomplish consistent style transfer from an arbitrary 2D style image to the 3D scene. Building on recent advances in 3D Gaussian Splatting (3DGS) [17], we optimize the 3DGS representation to generate stylized scenes with intricate artistic features and high visual quality. Our stylization process consists of two key steps: 1) We first recolor the 3D scene to align its color statistics with those of the style image. Simultaneously, we adopt a filter-based 3D Gaussian refinement scheme to minimize the impact of floaters in the reconstruction, which is crucial for ensuring the final stylization quality. 2) We then optimize 3D Gaussians further by exploiting nearest-neighbor feature matching style loss to capture detailed local style patterns. During the optimization, we incorporate a depth preservation loss and regularization terms to preserve the geometric features of the scene. Finally, we introduce how to achieve flexible control over the color, scale, and spatial attributes in our 3DGS stylization (Sec. 3.3). As a result, users can create stylized renderings with customized artistic expression, and explore the consistently stylized scene in a free-view manner. The pipeline of our method is shown in Fig. 2.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting [17] encapsulates 3D scene information using a set of 3D colored Gaussians. This technique exhibits rapid inference speeds and exceptional reconstruction quality compared to NeRF. To represent the scene, each Gaussian is described by a centroid $p=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D vector $s\in\mathbb{R}^{3}$ for scaling, and a quaternion $q\in\mathbb{R}^{4}$ for rotation. Additionally, an opacity value $\alpha\in\mathbb{R}$ and a color vector $c$ represented in the coefficients of a spherical harmonic (SH) function of degree 3 are used for fast alpha-blending during rendering. These trainable parameters are collectively symbolized by $G_{\theta_{i}}$, where $G_{\theta_{i}}=\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$, representing the parameters for the $i$-th Gaussian. To visualize the 3D Gaussians and supervise their optimization, 3DGS projects them onto the 2D image plane. The implementation leverages differentiable rendering and gradient-based optimization on each pixel for the involved Gaussians. The pixel color $c^{\alpha}$ is determined by blending the colors $c_{i}$ of those ordered Gaussians that overlap the pixel. This process can be formulated as:
@@ 3.2 Style Transfer to 3D Gaussian Splatting
Given a set of content images $\mathcal{I}_{content}$ that are captured from different viewpoints in the same scene, we first reconstruct its 3DGS model $G_{\theta}^{rec}$. Our goal is to transform $G_{\theta}^{rec}$ to a stylized 3DGS model $G_{\theta}^{sty}$ that matches the detailed style features of a given 2D style image $\mathcal{I}_{style}$ while maintaining the content of the original scene.
Color Transfer. We first transfer the color distribution of the style image to the 3DGS $G_{\theta}^{rec}$, enhancing the alignment of hues with the style image. We will introduce how we enable a flexible color control later. Drawing inspiration from [24], we employ a linear transformation of colors in RGB space, followed by a color histogram matching procedure between the style images and the views of the training set. Let ${p_{s}^{i}}$ represent the set of all pixels in the style image, and ${p_{c}^{i}}$ denote the set of all pixels in the content images to be recolored. We then solve the following linear transformation to align the mean ($\mu_{*}$) and covariance ($\Sigma_{*}$) of color distributions between the content image set and the style image:
$p_{c}^{re}=\mathbf{A}p_{c}+b,\,c^{re}=\mathbf{A}c+b$ (2)
$s.t.\,\mu_{p_{c}^{re}}=\mu_{p_{s}},\Sigma_{p_{c}^{re}}=\Sigma_{p_{s}}$
where $\mathbf{A}\in\mathbb{R}^{3\times 3}$ and $b\in\mathbb{R}^{3}$ are the solution to the linear transformation, $p_{c}^{re}$ is a recolored pixel of the content images. The color parameter $c$ of 3DGS $G_{\theta}^{rec}$ is transformed to the recolored version $c^{re}$. Please refer to the supplementary document for the mathematical derivation.
Filter-based Refinement. After the color transfer step, the color attributes $c^{re}$ of 3DGS and the recolored rendering images $\mathcal{I}_{content}^{re}$ are consistent with the palette of the style image. However, some floaters in the original 3DGS will also be colored during this process, significantly affecting the quality of the stylization, as shown in Fig. 13. These colored floaters often result from Gaussians with excessively large scales or improper opacity values. To address this issue, we design a filter-based refinement process to exclude these floaters while ensuring reconstruction quality before stylization. We use the recolored content images $\mathcal{I}_{content}^{re}$ as supervision to fine-tune the 3DGS with the recolored color parameter $c^{re}$ and filter the 3D Gaussian floater at every certain number of fine-tuning iterations. The filtering is performed based on the scale and opacity of the 3D Gaussians. Specifically, Gaussians whose sizes are in the top kor whose opacities are in the lowest kwill be filtered out. The threshold kremains constant during the fine-tuning process. By maintaining a fixed threshold, we ensure that the reconstructed scene does not change significantly, as the fine-tuning process is relatively short. Such a refinement step minimizes the impact on the overall scene reconstruction quality and enhances color correspondence, while the GS filter removes the floaters. During fine-tuning, we incorporate the reconstruction loss from [17]:
@@
Stylization. We then employ an intuitive and effective strategy to learn 3D Gaussian stylization by leveraging features extracted by a pre-trained convolutional neural network (e.g., VGG [30]). The 3D Gaussians are optimized with a set of loss functions between training views and the style image. To transfer detailed high-frequency style features from a 2D style image to a 3D scene, we exploit the nearest neighbor feature matching concept inspired by [13, 55], where we minimize the cosine distance between the feature map of rendered images and its nearest neighbor in the style feature map. We extract the VGG feature maps $F_{style}$, $F_{content}$, and $F_{render}$ for $\mathcal{I}_{style}$, $\mathcal{I}_{content}^{re}$, and the rendered images $\mathcal{I}_{render}$, respectively. The nearest neighbor feature match loss is formulated as:
@@
where, $(i*,j*)=\arg\min_{i^{\prime},j^{\prime}}D(F_{render}(i,j),F_{style}(i^{\prime},j^{\prime}))$, and
$D(a,b)$ is the cosine distance between two feature vectors $a$ and $b$. To preserve the original content structure during the stylization, we additionally minimize the mean squared distance between the content feature $F_{content}$ and the stylized feature $F_{render}$:
@@
where $H$ and $W$ represent the size of rendered images. While the content loss $\mathcal{L}_{content}$ mitigates the risk of excessive stylization, optimizations applied to the geometric parameters of 3D Gaussians can still induce changes in scene geometry. To address this, we introduce a depth preservation loss to maintain general geometric consistency without relying on any additional depth estimator. It should be noticed that there is a trade-off between geometric optimization and depth preservation loss, allowing our stylization to effectively transfer style patterns while avoiding potential damage to geometric content. Specifically, in the color transfer stage, we initially generate the original depth map $D_{origin}$ by employing the alpha-blending method of 3DGS. The alpha-blended depth $d^{\alpha}$ of each pixel is calculated as $d^{\alpha}=\sum_{i}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value for $i$-th Gaussian. During the stylization, we minimize the $L_{2}$ loss between the rendered depth map $D_{render}$ and the original one $D_{origin}$:
@@
We also add some regularization terms that perform on the changes of scale $\Delta s$ and opacity $\Delta\alpha$ in $G_{\theta}^{sty}$:
@@
Figure 3: Color Control Results. Our approach facilitates versatile color management in stylized outputs, allowing users to retain the scene’s original hues or apply distinct color schemes from alternative style images. Users can choose to transfer the entire style, only the pattern style, or a mix of arbitrary patterns and color styles.